OpenStudy (megannicole51):

what is the ratio of this sequence... 1+1/2+1/3+1/4+1/5.....

OpenStudy (megannicole51):

@Hero do u know how do find the ratio?

hero (hero):

Are you looking for the COMMON ratio?

OpenStudy (megannicole51):

the question just says give the first term, which is 1, and the ratio between successive terms

hero (hero):

Okay, so the common ratio is what they want

OpenStudy (megannicole51):

i believe so:)

OpenStudy (megannicole51):

and this is a geometric sequence correct?

hero (hero):

It appears that way. Hang on...

OpenStudy (megannicole51):

okay:)

hero (hero):

Well, I'm pretty sure the common ratio is \(\dfrac{1}{n + 1}\)

OpenStudy (megannicole51):

yeah i have that written down but isnt there more than i need?

hero (hero):

But I don't think that is geometric

OpenStudy (megannicole51):

what makes a sequence geometric?

hero (hero):

If it was geometric, then you could multiply by the common ratio to produce the next term.

OpenStudy (anonymous):

dont you think its a HP

OpenStudy (anonymous):

?

hero (hero):

In this case, we have the common ratio but we can't multiply it by the previous term to produce the next term.

hero (hero):

So this is neither arithmetic or geometric.

OpenStudy (megannicole51):

we cant just assume the next terms is (1/6) and that makes it a geometric sequence?

hero (hero):

You can assume what the next terms will be, but you cannot conclude that the sequence is geometric.

hero (hero):

The only thing you can do is try to come up with some kind of recurrence equation.

OpenStudy (megannicole51):

oooh okay....so like with 1-x+x^2-x^3+x^4....would this be geometric?

hero (hero):

Now we're skipping to something else completely different?

OpenStudy (megannicole51):

*wouldn't

OpenStudy (megannicole51):

well we already established that the other question isnt geometric and we already have the ratio...im just trying to understand/distinguish when a sequence is a geometric sequence and these are just some practice problems in my book.

hero (hero):

Not every sequence will be geometric or arithmetic.

hero (hero):

Here's a geometric sequence: 1, 2, 4, 8, 16, 32...

hero (hero):

It has a common ratio of 2

hero (hero):

It is geometric because you can \(\color\red{\text{multiply}}\) the previous term by 2 to get the next term.

hero (hero):

The key word is multiply

OpenStudy (megannicole51):

okay that makes sense...so then how would you find the common ratio when you have a + - + - in the sequence...ive been seeing a lot of problems like that in my book but there aren't any examples that show how to solve it or determine if it is a geometric sequence

hero (hero):

You'd have to post one of them first

hartnn (hartnn):

1-x+x^2-x^3+x^4...

OpenStudy (megannicole51):

that would work lol

OpenStudy (megannicole51):

do u want me to post it on a different thing?

hartnn (hartnn):

dive next term by current term, -x/1 = ... x^2/-x =... -x^3/x^2 =.... are all the answers same ? if yes, than its geometric with common ratio as the answer to those. if not, its not geometric

hartnn (hartnn):

and bdw, there is NO common ratio for initial sequence, 1,1/2,1/3,1/4 ... its an harmonic sequence.

OpenStudy (megannicole51):

-x right? and it is geometric and thank you:)

hartnn (hartnn):

correct. :)

OpenStudy (megannicole51):

yay!

hero (hero):

@hartnn nice way of showing it without giving the answer.

OpenStudy (megannicole51):

you both did a great job thank you so much:)

hero (hero):

yeah, if it were me, I would have just said -x because the way I approach these I look for a direct relationship between consecutive terms, but @hartnn demonstrated how to properly find a geometric relationship. If you divide the current term by the previous term, then you should be able to find a common ratio if the sequence is geometric.

OpenStudy (megannicole51):

yeah thats a good way to figure it out....and i can do that with every problem right?

hartnn (hartnn):

every problem of geometric sequence, yes. for arithmetic sequence, you would have to check the difference between the terms, next term - current term = constant = common difference d

OpenStudy (megannicole51):

oooh okay

hartnn (hartnn):

for harmonic sequence, take the reciprocal of each term and see whether they are in arithmetic progression, like your initial problem, reciprocals are,1,2,3,4,5.... common difference = 2-1 = 3-2 = 4-3 = .... - 1

hartnn (hartnn):

** = 1

OpenStudy (megannicole51):

omgsh these kinds of problems have so many steps and formulas and tricks to remember!

OpenStudy (anonymous):

theres no proper formula to get the sum of a HP.. you just have to solve it..

hartnn (hartnn):

there are things to understand too, have a look at these for details and other formulas : http://openstudy.com/users/hartnn#/updates/503bb2a0e4b007f9003103b0

OpenStudy (megannicole51):

awesome thank you!!!

hartnn (hartnn):

you're welcome ^_^

OpenStudy (megannicole51):

if i post another question can you help? it has to do with sums of series

hartnn (hartnn):

i will surely try :)